OpenStudy (anonymous):

I want to build a right triangular deck behind my house. The hypotenuse of the deck will be 25 feet long. The other two sides of the deck will be equal in length. What is the length of the deck? 17.68 feet 15.98 feet 10.34 feet 9.32 feet

OpenStudy (studygurl14):

Use the Pythagorean Theorem. \[a^2+b^2=c^2\] However, since we know the two sides of the deck are equal, it makes it easier. a =b, so... \[a^2+(a)^2 = c^2\] \[2a^2=c^2\] the hypotenuse is 25 \[2a^2=25\] Can you do the rest?

OpenStudy (anonymous):

i got 3.53... i think i did something wrong

OpenStudy (anonymous):

2a^2=25 divide both sides by two gets you a^2=12.5 and I dont know where to go after that @StudyGurl14

OpenStudy (anonymous):

@heavenfun13 can you help me?

OpenStudy (studygurl14):

take the square root of both sides

OpenStudy (studygurl14):

Oops, so sorry. This was actually my mistake. I didn't do 25^2

OpenStudy (studygurl14):

hold on...

OpenStudy (studygurl14):

\[2a^2=25^2\] Now solve for a

OpenStudy (anonymous):

17.677669529663688110021109052621

OpenStudy (studygurl14):

Great job, so the answer is A